OpenStudy (anonymous):

1/3(6x - 12) = 4(1/2x + 1) -2

OpenStudy (turingtest):

1\[{1\over3(6x - 12)} = 4({1\over2x} + 1) -2\]or 2\[{1\over3}(6x - 12) = 4({1\over2x} + 1) -2\]or 3\[{1\over3}(6x - 12) = 4({1\over2}x + 1) -2\]or none of the above? If it is above say number 1,2, or 3